# Preprint Manuscript

### v3

@@ 2.2 Roadmap
The restriction to $\mathbb{N}$ rather than $\mathbb{Z}$ is essential. Cubic equations over $\mathbb{Z}$ admit decidability through Thue equations (Baker’s transcendence methods), Hasse principle obstructions, and elliptic curve descent via Mordell–Weil theory. Undecidability over $\mathbb{N}$ exploits combinatorial complexity introduced by non-negativity constraints, enforced polynomially through guard-gadgets: replacing constraints $E(\mathbf{x})=0$ of degree at most two with the guarded system $u\cdot E(\mathbf{x})=0$, $u=1+v^{2}$ maintains cubic degree while forcing $u\geq 1$. Naive activation variables or difference-of-squares encodings produce degree-four terms, necessitating this construction.
Stage Configuration Implication
Step 1 $\mathcal{A}_{1}$ exists $\delta=2$ decidable (four-squares)
Step 2 $\mathcal{A}_{1}\land\mathcal{A}_{2}$ exist Can reduce any $\delta=6$ to equivalent $\delta=3$
Step 3a Assume $\mathcal{A}_{3}$ exists Contradiction: $\mathcal{A}_{1}\land\mathcal{A}_{2}\land\mathcal{A}_{3}$ decides PA
Step 3b Reject hypothesis Conclusion: $\delta=3$ undecidable over $\mathbb{N}$
Table 1: Logical structure of the impossibility proof.
The technical realization is a witness of a single cubic polynomial $P_{\text{hard}}(\mathbf{x})$ with $m=O(K^{3}N^{3})$ variables, where $N$ is the proof length and $K=O(\log(\max_{i}f_{i}))$ bounds Zeckendorf digit count. Every intermediate constraint—after guard-wrapping—has degree $\leq 3$ (by explicit bookkeeping, see 2); sum-of-squares merging yields degree $\leq 6$, reduced to cubic via monomial shielding (Definition 6.1) with $O(K^{3}N^{3})$ auxiliary variables
$P_{\text{merged}}=\sum_{i}P_{i}^{2}\in\delta(6)$ (2.3)
followed by recursive monomial shielding: each monomial $x^{a}y^{b}z^{c}$ with $a+b+c>3$ factors as $p\cdot q$ with $\delta(p),\delta(q)\leq 3$, introducing auxiliary $w$ via $w=p\cdot q$. Iteration terminates in $O(1)$ rounds, yielding a single cubic polynomial $P_{\text{hard}}(\mathbf{x})$ with $O(K^{3}N^{3})$ variables. The variable inflation from systems to single equations reflects the cost of degree-preserving aggregation.
Our proof by contradiction states: No universal cubic equation $P_{\text{univ}}(n,\mathbf{x})$ can encode all halting problems: such a polynomial would render Chaitin [1975] constant
$\Omega=\sum_{M\text{ halts}}2^{-|M|}$ (2.4)
computable by enumeration. Our construction avoids this through non-uniformity: $P_{\text{hard}}$ depends on the uncomputable proof length $N$. This applies equally to systems and single equations, though systems admit sparse structure while single equations become dense under sum-of-squares.
[…]
The cubic barrier represents a genuine phase transition in Diophantine complexity over $\mathbb{N}$, arising from arithmetic structure, non-negativity, and computational irreducibility: systems achieve cubic degree through encoding design, while single equations require degree-reduction post-processing.
Figure 1: A high-level overview of the reduction from provability in Peano Arithmetic to the solvability of cubic Diophantine systems. The construction proceeds through seven stages, each maintaining polynomial bounds on variable count and equation degree.
2.2.1 Encoding Proof Sequences (Stages 1–3)
A proof $\pi=\langle\varphi_{1},\ldots,\varphi_{N}\rangle$ is encoded via Gödel numbers $f_{i}\in\mathbb{N}$. The sequence is stored using two complementary representations:
@@ 2.2.1 Encoding Proof Sequences (Stages 1–3)
    $\beta$-Function Storage (Theorem 3.1): Existentially quantify over parameters $c,d\in\mathbb{N}$ satisfying
    $\beta(c,d,i)=f_{i}\quad\text{for all }i\in\{1,\ldots,N\},$ (2.5)
    where $\beta(c,d,i)$ denotes the remainder when dividing $c$ by $1+(i+1)d$. This is enforced via quadratic constraints (Lemma 4.1):
    $c$ $=q_{i}\cdot(1+(i+1)d)+r_{i},$ (2.6)
[…]
@@
    Zeckendorf Digits (Definition 3.3): Each $f_{i}$ is represented as a sum of non-consecutive Fibonacci numbers:
    $f_{i}=\sum_{\kappa=2}^{K}F_{\kappa}\cdot d_{i,\kappa},\quad d_{i,\kappa}\in\{0,1\},\quad d_{i,\kappa}\cdot d_{i,\kappa+1}=0.$ (2.8)
    This representation is unique and enables arithmetic verification without carry propagation (Exposition Exposition).
2.2.2 Proof Verification (Stage 4)
Each line $i$ must be justified either as an axiom or by modus ponens from earlier lines. This disjunction is encoded via boolean activation variables:
@@ 2.2.2 Proof Verification (Stage 4)
    Axiom Test (Lemma 4.5): For a finite axiom set $\{g_{1},\ldots,g_{M}\}$, the constraint
    $b_{ax,i}\cdot\sum_{\ell=1}^{M}(f_{i}-g_{\ell})^{2}=0,\quad b_{ax,i}\in\{0,1\}$ (2.9)
    forces $f_{i}\in\{g_{1},\ldots,g_{M}\}$ when $b_{ax,i}=1$ (since a sum of squares vanishes iff each term vanishes).

### v4

@@ 2.6 Constructive Reading
has computational content: it demands a realizer, i.e. an effective procedure which, given $(e,p,b)$, produces a code
$q=\mathsf{Close}(e,p,b)$ (2.22)
together with a correctness witness for (2.19). In other words, the closure principle is a uniform program-existence claim. This is precisely the sort of uniform witness-selection principle that realizability makes logically visible. One is not entitled to trade an external existential quantifier (“there exists some tt”) for an internal cubic solvability predicate without exhibiting a uniform method that performs this compilation.
Definition 2.15 (Uniform Cubic Closure Schema).
Let $\mathsf{Close}$ range over total computable functions (realizers). The uniform cubic closure principle is the schema:
$\exists\mathsf{Close}\;\forall e,p,b\;\Bigl(\mathsf{HaltsOut}(e,p,b)\ \leftrightarrow\ \mathsf{Sol}_{3}(\mathsf{Close}(e,p,b))\Bigr).$ (2.23)
Remark.
In classical logic, (2.23) merely asserts existence of a function with a correctness property. In the constructive layer, it is stronger: it asserts existence of a program implementing $\mathsf{Close}$. The subsequent refutation therefore targets the realizability content of (2.23), not merely its classical truth value.
Exposition.
If a uniform cubic closure operator existed, then unbounded halting-output predicates of the form
$\exists t\;\mathsf{Run}(e,p,t,b)$ (2.24)
could be uniformly reduced to cubic solvability queries. Since $\mathsf{Sol}_{3}$ is exactly the predicate decided by a hypothetical cubic solver, the combination
$\text{(cubic solver)}\ +\ \text{(cubic closure)}$ (2.25)
would yield a uniform decision procedure for a class of $\Sigma_{1}$ predicates sufficient to support diagonalization. Conceptually: bounded verification is decidable because the witness bound is fixed; closure converts “there exists a witness” into a single bounded-format question in $\mathcal{D}_{3}$. This is a global stabilization of $\Sigma_{1}$ existence in the cubic fragment.
Remark.
Gödelian non-uniformity is precisely the failure of uniform methods that decide or select witnesses for unbounded $\Sigma_{1}$ claims in general. The closure operator is such a witness-selection principle as it turns an unbounded existential over traces into a single cubic solvability instance, uniformly in parameters. Thus, at the constructive level, the closure principle is rejected.
Thesis 2.16.
There is no realizable (computable) function $\mathsf{Close}$ satisfying (2.23). Equivalently, the uniform cubic closure schema (2.23) admits contradiction in $\mathsf{HA}$.
Assume, for contradiction, that such a realizer $\mathsf{Close}$ exists. Then for each $(e,p,b)$ we can compute a cubic code $q=\mathsf{Close}(e,p,b)$ such that
$\exists t\;\mathsf{Run}(e,p,t,b)\ \leftrightarrow\ \mathsf{Sol}_{3}(q).$ (2.26)
If, in addition, we had any total correct solver for $\mathsf{Sol}_{3}$, we would obtain a uniform method deciding
$\exists t\,\mathsf{Run}(e,p,t,b)$ (2.27)
for the relevant class of
$(e,p,b),$ (2.28)
which suffices (by standard diagonalization) to collapse Gödelian non-uniformity. Since such a collapse is incompatible with Incompleteness for sufficiently strong effective theories, the assumption that a realizable $\mathsf{Close}$ exists must be rejected.
In particular, the failure is not algebraic but logical: $\mathsf{Close}$ would implement a uniform existential-elimination principle inside the cubic fragment. ∎
Remark.
The conclusion is intentionally stated at the level of realizability (program existence), not as an internal arithmetical theorem of $\mathsf{HA}$. The paper does not claim that $\mathsf{HA}$ proves $\neg\exists\mathsf{Close}$ as a closed formula. Rather, the claim is that (2.23) has no realizer.
2.7 Classical reflection and modus tollens
Exposition.
Layer III is a classical metatheoretic reading (e.g. in $\mathsf{PA}$) which treats the solver hypothesis as a definite arithmetical commitment. For fixed $e$, the sentence $\mathsf{Solve}_{3}(e)$ is classically determinate. The bounded encodings of Layer I are used only as representations of verification predicates, and the constructive analysis of Section 2.6 is used only to identify which uniform program-existence claims are untenable.
The reflection step required is minimal: we appeal to the soundness of the constructive reading for the relevant program-existence assertions (equivalently, to the metatheoretic correctness of the realizability interpretation for $\mathsf{HA}$ in this fragment). This allows us to pass from “no realizer exists” to the classical negation of the corresponding existence claim.
There is a useful formal analogue between (i) algebraic degree-preserving aggregation (Section 6.1) and (ii) the reflective step used here. Algebraically, maintaining $\delta\leq 3$ under aggregation forces arity growth: high-degree monomials are eliminated by introducing auxiliary variables via recursive shielding, replacing
$m=p\cdot q\;(\text{with }\delta(p),\delta(q)\leq 3)$ (2.29)
by a fresh variable $w$ constrained by $w=p\cdot q$. Iteration preserves degree but inflates the variable set.
Logically, maintaining the bounded $\Sigma_{1}$ format of verification predicates forces an analogous uniformity growth: bounded run instances
$\Phi_{e,p,t,b}=0$ (2.30)
represent $\mathsf{Run}(e,p,t,b)$ for each fixed $t$, but the unbounded existential closure
$\exists t\,\mathsf{Run}(e,p,t,b)$ (2.31)
can be internalized in the cubic fragment only by a uniform compilation principle (a closure operator) packaging the family $(\Phi_{e,p,t,b})_{t\in\mathbb{N}}$ into a single cubic instance. Thus, the same “cost” appears in two guises:
$\text{degree preservation}\ \leadsto\ \text{variable inflation}$ (2.32)
and
$\text{bounded verification}\ \leadsto\ \text{uniform closure demand}.$ (2.33)
In particular, any classical refutation of the uniform closure demand must quantify globally over all bounded instances; this is a compactness-like move at the metalevel, precisely mirroring the arity blow-up required by degree-preserving aggregation at the object level. Viewed informationally, such a global treatment amounts to extracting uniform content across all bounds, a pattern characteristic of incompleteness-by-complexity phenomena, cf. Chaitin [1975].
Proposition 2.17 (Solver Forces Closure).
For any $e$, if $\mathsf{Solve}_{3}(e)$ holds, then there exists a cubic closure operator $\mathsf{Close}$ satisfying (2.19) uniformly in $(e,p,b)$.
Remark.
Proposition 2.17 is the uniformity-extraction step: it isolates the closure principle as an implicit consequence of total correctness of a solver for $\mathcal{D}_{3}$.
Theorem 2.18 (No Total Correct Cubic Solver).
$\neg\exists e\;\mathsf{Solve}_{3}(e).$ (2.34)
By Modus Tollens, assume for contradiction that $\exists e\,\mathsf{Solve}_{3}(e)$ and fix such an $e$. By Proposition 2.17, $\mathsf{Solve}_{3}(e)$ entails the existence of a uniform cubic closure operator $\mathsf{Close}$ satisfying (2.19). But by Thesis 2.16, no realizable $\mathsf{Close}$ exists. By the reflection principle invoked above (soundness of the constructive reading for program-existence claims), we may therefore classically conclude $\neg\exists\mathsf{Close}$ with the required correctness property, contradicting the consequence of $\mathsf{Solve}_{3}(e)$. Hence $\neg\mathsf{Solve}_{3}(e)$ for all $e$, i.e. $\neg\exists e\,\mathsf{Solve}_{3}(e)$. ∎
Logic 2.19.
Hilbert’s Tenth Problem is a uniform decision problem. Theorem 2.18 negates precisely the global solver hypothesis for the cubic fragment $\mathcal{D}_{3}$. No claim is made that the refutation is formalizable in $\mathsf{HA}$; the argument is metatheoretic and uses a classical reflection step.
Equally, the argument does not rest on exhibiting a particular “mysterious” cubic equation, nor on constructing a universal cubic polynomial in the MRDP sense. The obstruction is logical: any solver would have to realize an unbounded existential-closure principle in the cubic regime, and this principle is shown to be inconsistent at the constructive level.
3 Preliminaries
3.1 encoding of finite sequences
Exposition.
The $\beta$-method stores a finite sequence $(a_{0},\dots,a_{n})$ inside a single integer $c$ by arranging congruences with pairwise coprime moduli. In our Diophantine systems we will not treat “Chinese Remainder” as a primitive operation; instead we encode it by the usual quotient–remainder equations plus an explicit bound.
Theorem 3.1 ($\beta$-encoding of a finite sequence).
Let $a_{0},a_{1},\ldots,a_{n}\in\mathbb{N}$ be given. There exist $c,d\in\mathbb{N}$ such that for every $i\in\{0,1,\ldots,n\}$ the remainder of dividing $c$ by
$m_{i}\;\equiv\;1+(i+1)d$ (3.35)
is exactly $a_{i}$. Equivalently, for each $i$ there exist $q_{i}\in\mathbb{N}$ with
$c\;=\;q_{i}\,m_{i}\;+\;a_{i}\quad\text{and}\quad 0\leq a_{i}<m_{i}.$ (3.36)
Choose $d$ large enough that $d>\max_{i}a_{i}$ and divisible by $\mathrm{lcm}(1,2,\ldots,n+1)$. Then $a_{i}<m_{i}$ for all $i$, and the moduli $m_{i}=1+(i+1)d$ are pairwise coprime. Indeed, if a prime $p$ divides $m_{i}$ and $m_{j}$ then it divides their difference (ji)d(j-i)d. If $p\mid d$ then $m_{i}\equiv 1\pmod{p}$, impossible; hence $p\nmid d$, so $p\mid(j-i)\leq n$, but then $p\mid\mathrm{lcm}(1,\ldots,n+1)\mid d$, contradiction. By the Chinese remainder theorem, there exists $c$ with $c\equiv a_{i}\pmod{m_{i}}$ for all $i$, which yields (4.64). ∎
Remark.
In the construction later, the values $a_{i}$ are themselves variables (they will be the Gödel Numbers $f_{i}$ of proof lines). Thus we implement $\beta(c,d,i)=f_{i}$” by introducing quotient variables $q_{i}$ and enforcing
$c=q_{i}(1+(i+1)d)+f_{i}\quad\text{together with}\quad 0\leq f_{i}<1+(i+1)d.$ (3.37)
The only non-equational ingredient is the inequality, which we convert to equations using the four-squares lemma below.
3.2 Bounds as equations
Lemma 3.2 (Four-squares bound encoding).
For $r,m\in\mathbb{N}$ the constraint $0\leq r<m$ is equivalent to the existence of
$a_{1},a_{2},a_{3},a_{4},b_{1},b_{2},b_{3},b_{4}\in\mathbb{N}$ (3.38)
such that
$r$ $=a_{1}^{2}+a_{2}^{2}+a_{3}^{2}+a_{4}^{2},$ (3.39)
$m-1-r$ $=b_{1}^{2}+b_{2}^{2}+b_{3}^{2}+b_{4}^{2}.$ (3.40)
In particular, the bound $0\leq r<m$ can be enforced using polynomial equations of degree $2$.
If (4.65) holds then $r\geq 0$. If (4.66) holds then $m-1-r\geq 0$, hence $r\leq m-1$, i.e. $r<m$. Conversely, if $0\leq r<m$ then both $r$ and m1rm-1-r are nonnegative integers, hence each is a sum of four squares by Lagrange’s Theorem, cf. Lagrange [1770]. ∎
Remark.
For later completeness arguments it is useful that, given numerical $r$ and $m$, one can effectively find witnesses for (4.65)–(4.66). We only use existence and computability; we do not rely on fine-grained complexity estimates here.
3.3 Zeckendorf digits
Definition 3.3 (Zeckendorf representation).
Let $F_{1}=F_{2}=1$ and $F_{k+2}=F_{k+1}+F_{k}$. Every $n\in\mathbb{N}_{>0}$ has a unique representation
$n=\sum_{\kappa\in S}F_{\kappa},$ (3.41)
where $S\subseteq\{2,3,4,\ldots\}$ contains no consecutive indices.
Lemma 3.4 (Polynomial encoding of Zeckendorf digits).
Fix a cutoff $K\geq 2$. To express that $n$ is represented using indices $\leq K$, introduce variables $d_{2},\ldots,d_{K}\in\mathbb{N}$ and impose:
@@ 3.3 Zeckendorf digits
    $d_{\kappa}-d_{\kappa}^{2}=0$ for all $\kappa\in\{2,\ldots,K\}$,
@@
    $d_{\kappa}d_{\kappa+1}=0$ for all $\kappa\in\{2,\ldots,K-1\}$,
  3. 3.
    $n=\sum_{\kappa=2}^{K}F_{\kappa}d_{\kappa}$.
All constraints have degree $\leq 2$.
Over $\mathbb{N}$, $d_{\kappa}-d_{\kappa}^{2}=0$ forces $d_{\kappa}\in\{0,1\}$. The products enforce non-adjacency. The reconstruction is linear in the digits. ∎
Remark.
In base-$p$ digit encodings, verifying addition typically requires explicit carry variables and higher-degree constraints. With Zeckendorf Digits, we instead enforce the value equality (e.g. $f_{i}=f_{j}+f_{k}$) and separately require that $f_{i}$ has a valid Zeckendorf Digit vector. Uniqueness of Zeckendorf Representation then pins down the digits of the sum without introducing a carry-propagation gadget.
3.4 Carryless pairing
Definition 3.5 (Support, rank, delimiter).
Let $n\in\mathbb{N}$ have Zeckendorf Digits $(d_{\kappa})_{\kappa\geq 2}$ as in Lemma 3.4, and define its support
$Z(n)\;\equiv\;\{\kappa\geq 2:d_{\kappa}=1\}.$ (3.42)
Define the rank $r(n)$ by
$r(n)\;\equiv\;1+\max\bigl(Z(n)\cup\{1\}\bigr),$ (3.43)
so $Z(n)\subseteq\{2,3,\ldots,r(n)-1\}$. Define the delimiter $B(n)\equiv 2r(n)$.
Definition 3.6 (Typed carriers).
For $x\in\mathbb{N}$ define the finite carriers
$\tau_{x}=\{\,e\in\mathbb{N}:e<r(x)\,\},\quad\varepsilon_{x}=\{\,2e:e\in\tau_{x}\,\},$ (3.44)
and for $y\in\mathbb{N}$ the odd band
$\omega_{x,y}=\{\,B(x)+(2j-1):j<r(y)\,\}.$ (3.45)
Lemma 3.7 (Typed finiteness).
For each $x,y\in\mathbb{N}$ the sets $\tau_{x}$, $\varepsilon_{x}$, and $\omega_{x,y}$ are finite. Their bounds depend only on $r(x)$ and $r(y)$, fixed from the Zeckendorf Supports of $x$ and $y$.
Immediate from the definitions: each set is cut off by $r(x)$ and/or $r(y)$. ∎
Definition 3.8 (Typed carryless pairing).
The Typed Carryless Pairing defined as
$\pi_{\mathrm{CL}}^{\tau}(x,y)=\!\!\!\!\!\sum_{e\in Z(x)\subseteq\tau_{x}}\!\!\!\!\!F_{2e}\;+\;\!\!\!\!\!\sum_{j\in Z(y)\subseteq\tau_{y}}\!\!\!\!\!F_{\,B(x)+(2j-1)}.$ (3.46)
The even band encodes $x$ and the odd band encodes $y$, separated by the delimiter $B(x)=2r(x)$ of Definition 3.5.
Exposition.
The carriers supply the finite index domains across which sums, tests, and bounded recursions occur. In particular, once $x$ and $y$ are fixed, all indices used in $\pi_{\mathrm{CL}}^{\tau}(x,y)$ lie in the finite set
$\varepsilon_{x}\;\cup\;\omega_{x,y},$ (3.47)
and no bound depends on intermediate values discovered during evaluation. This is the formal sense in which the pairing is “carryless”: the encoding uses disjoint digit bands rather than arithmetic carry propagation.
3.5 Combined encoding
Definition 3.9 (Guarded Constraint).
$u\cdot E(\mathbf{x})$ $=0,$ (3.48)
$u-1-v^{2}$ $=0.$ (3.49)
Proposition 3.10 (Guard Semantics).
The guarded constraint enforces:
@@ 3.5 Combined encoding
    $u=1+v^{2}\geq 1$ for all solutions (from equation (3.49)),